If S= {e*, e*, xe* is a fundamental solution set for the homogenous differential equation with constant coefficients dy a2 dy +a dx2 d y + a y 0, dx dx find the values of a ,i= 0, 1, 2, 3 (5 markah/marks)
Dengan nilai yang diperolehi dari (a), selesaikanlah masalah nilai awal yang berikut: (b) day 2 (1+2e) d2 y a a dx dy ax a ax dengan 0)1,y' (0)- 2,y (0)-3. By using the values obtained in (a), solve the initial value problem: d2 y dy ay 2 (1+2e ) d'y a d x * ax ax with 0)1,y' (0) 2,y (0) =3.
Answer #1

+a,9=0. D-t a, D+0, D a. Le, ee Solution i Rots of the equedi0, is -than 40= e It Cast%)- PpM Saluhion dat) b Capariny atTon= a (1t e) Gete (de -D+1 DD-Dtl DD-Dt 0D-Ot 2e denonun became S D2-D-Dt! 2x again denotunatov 6 D- Zre. (+ ae = 2+xe* 2 X = eyo=+G G+Gm+x¢c* a=eestaJ =C st2+2] +Fst+srju#j = Gte acstta ++ca) n+n°| Subhtate akre vaseg in y, y, ytu CrtCLETI yo)= 2tC